Consider the combustion of 1-propanol (C3H8O):
Calculate the mass of carbon dioxide produced from 100 kg of 1-propanol.
Calculate the mass of oxygen consumed in the reaction above.
1
Expert's answer
2020-03-10T09:08:16-0400

"C_3H_8O+4.5O_2\to3CO_2+4H_2O"

"60\ g" Propanol produce "44\times 3=132\ g" of "CO_2"

"100 \ Kg" Propanol will produce "\frac{132}{60}\times100\ Kg =220\ Kg" of "CO_2"


"60\ g" Propanol consume "4.5\times 32=144 \ g" of "O_2"

"100 \ Kg" Propanol consume "\frac{144}{60}\times 100 \ Kg=240\ Kg" of "O_2" .